OpenStudy (anonymous):

Given the following function, find f (− 3 x) . f ( x )  =  − 8 x 2 − 6

OpenStudy (sidsiddhartha):

\[f(x)=-8x^2-6\] so to evaluate f(-3x) just put -3x in place of x \[f(-3x)=-8(-3x)^2-6=-72x^2-6\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

if that is true wouldn't it cancel out the negatives \[f(-3)=-8(-3x)^2-6=72x^2-6\]

OpenStudy (sidsiddhartha):

-8*(-3x)^2=(-8)*(-3x)*(-3x) so one negetive remains

OpenStudy (anonymous):

got it

OpenStudy (sidsiddhartha):

:)